# Return-Aligned Decision Transformer

### v5

@@ A.2 Additional Ablation Study
We conduct an ablation study on the adaptive scaling of SeqRA in the MuJoCo domain. The experimental setup is the same as described in Sec.5.5. The results are summarized in Tab.5. These results demonstrate that introducing adaptive scaling (RADT w/o StepRA) outperforms the case where only the attention mechanism of SeqRA is used (RADT w/o StepRA and Adaptive Scaling). This suggests that adaptive scaling effectively merges $z$, which includes information from the return sequence, with the state-action sequence $\tau_{sa}$.
Table 5: Absolute $\downarrow$ comparison in the ablation study on the adaptive scaling in SeqRA. The way to interpret this table is the same as that of Tab 3.
@@
We conduct an ablation study for SeqRA and StepRA in the Atari domain. The experimental setup is the same as in Sec. 5.5. The results are summarized in Tab. 6. These results indicate that introducing either SeqRA or StepRA independently is effective. Furthermore, combining both aligners results in the minimum absolute error, implying a complementary relationship between SeqRA and StepRA. RADT w/o StepRA has the same or superior performance compared to RADT w/o SeqRA. This suggests that SeqRA is advantageous for the long-term credit assignments required in the Atari domain.
@@ C.2 Hyperparameters
Number of layers $2$, hopper-m

### v6

@@ A.2 Additional Ablation Study
We conduct an ablation study on the adaptive scaling in SeqRA on the medium-replay dataset in MuJoCo and the Atari dataset. The experimental setup is the same as described in Sec.5.5. The results are summarized in Tab. 8 and Tab. 9. These results demonstrate that introducing adaptive scaling (RADT w/o StepRA) outperforms the case where only the attention mechanism of SeqRA is used (RADT w/o StepRA and Adaptive Scaling). This suggests that adaptive scaling effectively merges $z$, which includes information from the return sequence, with the state-action sequence $\tau_{sa}$.
Table 8: Absolute $\downarrow$ comparison in the ablation study for the adaptive scaling in SeqRA on the medium-replay dataset in MuJoCo domain. The way to interpret this table is the same as that of Tab 3.
@@
Table 9: Absolute $\downarrow$ comparison in the ablation study for the adaptive scaling in SeqRA on the Atari domain. The way to interpret this table is the same as that of Tab 3.
@@
A.3 Additional Analysis of Parameter Size
We compare RADT and models where we increase the number of blocks in DT to bring the number of parameters closer to RADT. Table 10 shows the parameters for each model on the medium-replay dataset of MuJoCo. Because RADT adds both SeqRA and StepRA to the transformer block, it has 1.6 times the number of parameters compared to DT. By increasing the number of blocks in DT to six, it surpasses the number of parameters of RADT. Therefore, we compare RADT with a DT that has six blocks and present the results on the medium-replay dataset of MuJoCo in Tab. RADT outperforms the DT with six blocks. On the other hand, we find that DT with 6 blocks performs worse than that with 3 blocks in the hopper and walker2d environments. This result suggests that simply increasing the number of parameters of DT does not improve return alignment, and that the improved model structure of RADT is effective.
Table 10: Comparison of parameters. Since the dimensions of the state and action spaces vary for each dataset, the model size differs depending on the dataset.
Environment DT(3 blocks) DT(5 blocks) DT(6 blocks) RADT(3 blocks)
ant 740,232 1,136,776 1,335,048 1,219,848
halfcheetah 727,686 1,124,230 1,322,502 1,207,302
hopper 726,147 1,122,691 1,320,963 1,205,763
walker2d 727,686 1,124,230 1,322,502 1,207,302
Table 11: Absolute $\downarrow$ comparison between RADT and DT with increased parameters. We use the medium-replay dataset in MuJoCo. The way to interpret this table is the same as that of Tab 1.
Environment DT(3 blocks) DT(6 blocks) RADT(3 blocks)
ant $23.2\pm 1.3$ $16.8\pm 4.5$ $\mathbf{3.5}\pm 0.5$
halfcheetah $7.2\pm 1.8$ $5.5\pm 0.5$ $\mathbf{1.8}\pm 0.6$
hopper $12.2\pm 4.3$ $12.8\pm 2.4$ $\mathbf{4.4}\pm 0.5$
walker2d $8.1\pm 0.3$ $10.0\pm 0.7$ $\mathbf{4.0}\pm 0.8$
@@ C.2 Hyperparameters
Number of blocks $2$, hopper-m
@@
Appendix D Additional Real-World Application Example
We describe the additional real-world application that requires return alignment, as mentioned in Sec. 1.
Example 3 (Agent-based modeling and simulation).